# The causal structure of galactic astrophysics

### v1

@@ 4.2 Real data
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift influences the apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance. Mass is seen to causally determine size (rather than vice versa, as would be possible given simply the mass–size relation), suggesting inside-out growth of discs and size expansion via mergers. It also determines Sérsic index, agreeing with the idea that bulge growth and morphological transformation are primarily consequences of hierarchical mass assembly. The absence of link from morphology to mass disfavours simplistic models where concentration alone sets stellar mass. The link from SFR to absolute magnitude reflects the brightening of galaxies in optical bands due to recent star formation.
The uncertain edges between star formation, stellar mass, and morphology highlight the complexity of baryonic processes. The ambiguous edge between stellar mass and luminosity is unsurprising, since mass estimates are derived from photometry and strongly depend on mass-to-light ratios. The circle endpoints highlight the possible role of latent factors not included in the analysis—such as dust attenuation, gas content, and halo environment—which can drive correlations and obscure true directions. The graph does not unambiguously support a picture in which star formation determines morphology on short timescales, or that mass quenching is the sole pathway. It does however imply that the backbone of galaxy evolution—mass driving size and morphology, and star formation driving luminosity—is recoverable directly from survey data. This is highly promising for future, more sophisticated applications of the methodology.
Figure 3: The PAG of the NSA data. Confident causal structures are indicated by directed edges, while less confident associations (circle endpoints) may be impacted by latent confounders.
@@ 5 Discussion and Conclusion
The application of causal discovery to astrophysics is largely virgin territory. By enabling the directions of physical links to be established, it provides a significant information overlay on (even machine learning-based) correlation analyses, helping to constrain theories which postulate the physical mechanisms governing the data. Such theories essentially correspond to DAGs, so causal discovery can be seen as a method for inferring theories (as far as is possible) directly from data.
@@
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 3. This supports a hierarchical and mass-driven framework of galaxy evolution while indicating the complexities in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics.
@@
In the near term there are several ways in which this analysis could be extended. First, many of the causal links in Fig. 3 reflect observational or selection effects rather than physical mechanisms. The data could be finessed to minimise these, for example by conditioning properties on redshift or constructing combinations of variables less prone to selection biases or trivial correlations. Second, the several ambiguous (circle) endpoints indicate the potential impact of latent variables not included in the dataset. By folding in such properties as gas mass, metallicity, dust attenuation and environmental density these ambiguities could be resolved, providing a clearer picture of the overall flow of causality. There is of course a huge range of further data across astrophysics that could profitably be interpreted through a causal discovery lens.
@@
There is room for improvement on the theoretical side too. While we showed good performance, our method still relies on choices of conditional independence tests and scoring which have not been explored exhaustively. This could be investigated by applying causal discovery to cosmological simulations, which have known physical mechanisms but more accurately capture likely astrophysical correlations. This would also provide a platform for investigating selection effects in detail, as well as revealing more clearly the causal graphs associated with candidate physical models. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity or non-Gaussianity), as has already been done when assuming causal sufficiency (e.g. Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2023; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery.
@@
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code and all other data will be made available on reasonable request to the corresponding author.

### v2

@@ 4.2 Real data
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift influences the apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance. Mass is seen to causally determine size (rather than vice versa, as would be possible given simply the mass–size relation), suggesting inside-out growth of discs and size expansion via mergers. It also determines Sérsic index, agreeing with the idea that bulge growth and morphological transformation are primarily consequences of hierarchical mass assembly. The absence of a link from morphology to mass disfavours simplistic models where concentration alone sets stellar mass. The link from SFR to absolute magnitude reflects the brightening of galaxies in optical bands due to recent star formation.
The uncertain edges between star formation, stellar mass and morphology highlight the complexity of baryonic processes. Several appear with circle endpoints, indicating algorithmic uncertainty about direction or the influence of confounding variables. The ambiguous link between stellar mass and luminosity is unsurprising, since mass estimates are derived from photometry and strongly depend on mass-to-light ratios. The circle endpoints highlight the possible role of latent factors not included in the analysis—such as dust attenuation, gas content, and halo environment—which can drive correlations and obscure true directions. The graph does not unambiguously support a picture in which star formation determines morphology on short timescales, or that mass quenching is the sole pathway. Instead, it suggests an intertwined system in which stellar mass remains the fundamental regulator, while star formation and morphology are mutually influenced by common drivers such as gas inflow and environment. Disentangling these physical drivers from observational and selection-induced structure (and external latent variables) requires further work, as indicated directly by the circle edges. The result does however imply that the backbone of galaxy evolution—mass driving size and morphology, and star formation driving luminosity—is recoverable directly from survey data. This is highly promising for future, more sophisticated applications of the methodology.
Figure 3: The PAG of the NSA data. Each node contains a colloquial parameter name in bold as well as the technical variable name in the NSA. Confident causal structures are indicated by directed edges, while less confident associations (circle endpoints) may be impacted by latent confounders.
@@ 5 Discussion and Conclusion
The application of causal discovery to astrophysics is largely virgin territory. By enabling the directions of physical links to be established, it provides a significant information overlay on (even machine learning-based) correlation analyses, helping to constrain theories that postulate the physical mechanisms governing the data. Such theories essentially correspond to DAGs, so causal discovery can be seen as a method for inferring theories (as far as is possible) directly from data.
@@
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 3. This supports a hierarchical and mass-driven framework of galaxy evolution while indicating the complexities involved in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics. Such methods are sometimes undervalued in favour of intervention-based methods involving randomised control trials.
@@
In the near term there are several ways in which this analysis could be extended. First, many of the causal links in Fig. 3 reflect observational or selection effects rather than physical mechanisms. The data could be refined to minimise these, for example by conditioning properties on redshift or constructing combinations of variables less prone to selection biases or trivial correlations. Second, the several ambiguous (circle) endpoints indicate the potential impact of latent variables not included in the dataset. By folding in such properties as gas mass, metallicity, dust attenuation and environmental density these ambiguities could be resolved, providing a clearer picture of the overall flow of causality. There is of course a huge range of further data across astrophysics that could profitably be interpreted through a causal discovery lens.
@@
There is room for improvement on the theoretical side too. While we showed good performance, our method still relies on choices of conditional independence tests and scoring which have not been explored exhaustively. Mismatches between the mock and real data could lead to errors being larger in the latter. Besides directly investigating other hyperparameter choices, mock datasets and algorithms, this could be investigated by applying causal discovery to cosmological simulations, which have known physical mechanisms but more accurately capture correlations likely to be present in astrophysical data. This would also provide a platform for investigating the loss of information due to selection effects, reveal more clearly the causal graphs associated with candidate physical models and calibrate how reliably different orientations can be distinguished. In parallel, bootstrap and stability analyses on the observational data could determine which edges are most robust.
General and efficient causal discovery methods remain under active development. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity, non-Gaussianity or heteroskedasticity), as has already been done when assuming causal sufficiency in the linear, non-Gaussian regime (e.g. Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2023; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery. Taken together, these steps would provide a more complete and reliable causal map, ultimately allowing causal discovery to move beyond recovering broad backbones to testing detailed models of how galaxies grow and evolve.
@@
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code used in this paper will be made publicly available on GitHub upon publication.